### v1

@@ 3.4 Query Group Attention
In our multi-view 2D decoder, we use the projected object center as the reference point for a 2D box. In the case of object truncation, we rely on the center of the bounding rectangle of the projected anchor. Finally, we utilize a shared 2D head across cameras to make predictions for the 2D bounding box and class.
@@ 3.6 Loss Functions
The training loss includes both 2D object detection loss and 3D detection loss: $\mathcal{L}=\mathcal{L}_{2d}+\mathcal{L}_{3d}$. We employ the same 3D detection loss in [37, 20]. For the 2D detection loss, we adopt $\mathcal{L}_{detr2d}$ from DETR-like methods [28, 45] and incorporate the observation angle [27] loss of a 3D box in the image view. Typically, we encode alpha-angle with sin and cos functions: $\mathcal{L}_{alpha}=\frac{1}{M}\sum_{i=1}^{M}|\sin(\theta_{i})-\hat{\sin}(\theta_{i})|+|\cos(\theta_{i})-\hat{\cos}(\theta_{i})|$, where $\theta_{i}$ is the observation angle of an object. And the 2D detection loss is extended to $\mathcal{L}_{2d}=\mathcal{L}_{detr2d}+\lambda_{alpha}\mathcal{L}_{alpha}$. $\lambda_{alpha}=0.5$ is used in the experiments.
@@ 4.2 Implementation Details
Table 1: Comparison results of 3D detection on nuScenes validation dataset. †The backbone benefits from perspective pertaining. $\Diamond$ indicates that we re-ran the experiments using the same setting of SimPB.
@@
Sparse4Dv3$\Diamond$ [20] ResNet101 $1408\times 512$ 0.520 0.614 0.499 0.262 0.309 0.195 0.197
SimPB† ResNet101 $1408\times 512$ 0.531 0.622 0.509 0.265 0.274 0.191 0.195
@@ 4.3 Main Results
3D Object Detection Performance. We compare SimPB with previous state-of-the-art methods on the validation split of nuScenes in Tab. 1. With a model input resolution of $704\times 256$ and ResNet50 as the backbone, SimPB surpasses previous state-of-the-art methods. For example, it outperforms Sparse4Dv3 [20] by 0.6% in mAP and 2.0% in NDS, also showing significant enhancements of 12.1% in mAOE. By using perspective pretrained backbone on nuImages [2], SimPB achieves a new record on the validate split, further improving the performance of DynamicBEV [42] by 2.3% in mAP and 2% in NDS. SimPB also achieves competitive performance by leveraging a larger ResNet101 backbone and a model input resolution of $1408\times 512$. It outperforms our re-run of Sparse4Dv3 with additional improvements of 1.1% in mAP and 0.8% in NDS.
@@ 4.4 Ablation Study
Table 5: Ablation studies on query allocation mechanism.
@@
Table 5: Ablation studies on query allocation mechanism.
@@
Table 7: The ablation studies of camera observation loss.
@@
Table 7: The ablation studies of camera observation loss.
@@ 0.C.3 Effect of Association
$\scriptsize\Phi(P_{3d}^{i},G_{2d}^{j})=\begin{cases}1&\text{if}~{}\text{Dist}(P_{3d}^{i},G_{3d}^{j})\leq\tau_{dis}~{}\&~{}\text{IoU}(P_{3d\rightarrow 2d}^{i},G_{2d}^{j})\leq\tau_{iou}~{}\&~{}\text{Cls}(P_{3d}^{i},G_{3d}^{j})=1\\ 0&\text{otherwise},\end{cases}$ (10)
@@
$\scriptsize\Psi(P_{3d}^{i},P_{2d}^{k})=\begin{cases}1&\text{if}~{}\Phi(P_{3d}^{i},G_{2d}^{j})=1~{}\&~{}\text{IoU}(P_{2d}^{k},G_{2d}^{j})\leq\tau_{iou}~{}\&~{}\text{Cls}(P_{2d}^{k},G_{2d}^{j})=1\\ 0&\text{otherwise}.\end{cases}$ (11)

### v2

@@ 3.4 Query Group Attention
In our multi-view 2D decoder, we use the projected object center as the reference point for a 2D box. In the case of object truncation, we rely on the center of the bounding rectangle of the projected anchor. Then, we utilize a shared 2D head across cameras to make predictions for the 2D bounding box and class. Finally, these predictions will be matched with the groundtruth in each groups through standard Hungarian matching algorithm used in DETR-like[28, 45, 22] methods.
@@ 3.6 Loss Functions
The training loss includes both 2D object detection loss and 3D detection loss: $\mathcal{L}=\mathcal{L}_{2d}+\mathcal{L}_{3d}$. Similar to [11, 15, 20], our method employ the same 3D detection loss with the depth maps as auxiliary supervision. For the 2D detection loss, we adopt $\mathcal{L}_{detr2d}$ from DETR-like methods [28, 45] and incorporate the observation angle [27] loss of a 3D box in the image view. Typically, we encode alpha-angle with sin and cos functions: $\mathcal{L}_{alpha}=\frac{1}{M}\sum_{i=1}^{M}|\sin(\theta_{i})-\hat{\sin}(\theta_{i})|+|\cos(\theta_{i})-\hat{\cos}(\theta_{i})|$, where $\theta_{i}$ is the observation angle of an object. And the 2D detection loss is extended to $\mathcal{L}_{2d}=\mathcal{L}_{detr2d}+\lambda_{alpha}\mathcal{L}_{alpha}$, where $\lambda_{alpha}$ is set to 0.5 in the experiments.
@@ 4.2 Implementation Details
Table 1: Comparison results of 3D detection on nuScenes validation dataset. †The backbone benefits from perspective pertaining.
@@
SimPB† ResNet101 $1408\times 512$ 0.539 0.629 0.475 0.260 0.280 0.192 0.197
@@ 4.3 Main Results
3D Object Detection Performance. We compare SimPB with previous state-of-the-art methods on the validation split of nuScenes in Tab. 1. With a model input resolution of $704\times 256$ and ResNet50 as the backbone, SimPB surpasses previous state-of-the-art methods. For example, it outperforms Sparse4Dv3 [20] by 0.6% in mAP and 2.0% in NDS, also showing significant enhancements of 12.1% in mAOE. By using perspective pretrained backbone on nuImages [2], SimPB achieves a new record on the validate split, further improving the performance of DynamicBEV [42] by 2.3% in mAP and 2% in NDS. SimPB also achieves competitive performance by leveraging a larger ResNet101 backbone and a model input resolution of $1408\times 512$. It outperforms Sparse4Dv3 with additional improvements of 0.2% in mAP and 0.6% in NDS.
@@ 4.4 Ablation Study
[…]
@@
Table 5: Ablation studies on query allocation mechanism.
[…]
@@
[…]
@@
Table 7: The ablation studies of camera observation loss.
[…]
@@ 0.C.3 Effect of Association
$\scriptsize\Phi(P_{3d}^{i},G_{2d}^{j})=\begin{cases}1&\text{if}~{}\text{Dist}(P_{3d}^{i},G_{3d}^{j})\leq\tau_{dis}~{}\&~{}\text{IoU}(P_{3d\rightarrow 2d}^{i},G_{2d}^{j})\geq\tau_{iou}~{}\&~{}\text{Cls}(P_{3d}^{i},G_{3d}^{j})=1\\ 0&\text{otherwise},\end{cases}$ (10)
@@
$\scriptsize\Psi(P_{3d}^{i},P_{2d}^{k})=\begin{cases}1&\text{if}~{}\Phi(P_{3d}^{i},G_{2d}^{j})=1~{}\&~{}\text{IoU}(P_{2d}^{k},G_{2d}^{j})\geq\tau_{iou}~{}\&~{}\text{Cls}(P_{2d}^{k},G_{2d}^{j})=1\\ 0&\text{otherwise}.\end{cases}$ (11)